OpenStudy (anonymous):

f(x) = 1/(x + 7) Find the limit. lim Δx→0 (f(x + Δx) − f(x))/Δx

OpenStudy (anonymous):

Ive tried a couple of different ways but i think im just multiplying the top and the bottom by the wrong values

OpenStudy (anonymous):

i was trying (x+\[Deltax\] + 7)(x+7)

OpenStudy (anonymous):

\[\frac{ d }{ dx}\frac{ 1 }{ x+7}\]=\[\lim_{\Delta x \rightarrow 0}\frac{ \frac{ 1 }{ x+\Delta x+7 }-\frac{ 1 }{ x+7 } }{ \Delta x}\]=\[\lim_{\Delta x \rightarrow 0}\frac{ -1}{ (x+7)(x+\Delta x+7) }\]

OpenStudy (anonymous):

hmmm how does that help?

OpenStudy (anonymous):

and how exactly did you get to that point?

OpenStudy (anonymous):

@YadielG are you want to differentiate this f(x) from definition of differentiation?

OpenStudy (anonymous):

I know that the second equation is how you define a derivative using limits, but i just want to find the value of the limit using that second equation.

OpenStudy (anonymous):

Im still new to Calculus 1 so im not too knowledgeable about the terminology and processes yet

OpenStudy (anonymous):

i know the equation does have a limit that exists and that it is not zero, but im just having trouble with the algebra getting to that value

OpenStudy (anonymous):

you can find the value of limit, just substitute Delta x=0 in this

OpenStudy (anonymous):

http://prntscr.com/8ac5if

OpenStudy (anonymous):

well that gives you undefined because of the delta x on the bottom

OpenStudy (lynfran):

i like working wih h instead of delta x ... \[\lim_{h \rightarrow 0}\frac{ (\frac{ 1 }{ x+7 }+h)-(\frac{ 1 }{ x+7 }) }{ h }\]\[\lim_{h \rightarrow 0}\frac{ \frac{ 1+hx+7h-1}{ (x+7) } }{ h }\]\[\lim_{h \rightarrow 0}\frac{ h(x+7) }{ (x+7) }*\frac{ 1 }{ h }\]

OpenStudy (anonymous):

i thought the goal in this problem was to get rid of all of the x's and just have delta x in order to find the limit

OpenStudy (anonymous):

fist must simply it before substitution by zero.

OpenStudy (anonymous):

@LynFran how can you pull out h from the bottom of the small fraction like that?

OpenStudy (anonymous):

@ASAAD123 Yes i understand that much but what im really having trouble with is simplifying it

OpenStudy (dinamix):

u sould the answer f'(0)= -1/49

OpenStudy (anonymous):

@dinamix yeah it's easy if i just take the derivative but i need to do it using that limit equation

OpenStudy (lynfran):

by finding its common denominator

OpenStudy (anonymous):

@ASAAD123 yepp i got to that point

OpenStudy (lynfran):

shouldnt it be....??